# Bayesian Perspective for Orientation Estimation in Cryo-EM and Cryo-ET

### v1

@@ 4.2 Empirical results for volume reconstruction and the “Einstein from Noise” phenomenon
A few observations can be made from both Figures 4 and 5. First, in the case of high SNR (i.e., as $\sigma\to 0$), the volume reconstruction is similar whether using the MAP estimator or the MMSE estimator. This similarity is theoretically inferred by Proposition 2.1. However, as the SNR decreases, the reconstructions diverge, with the volumes reconstructed using the MMSE estimator showing a better correlation with the true volume. Second, in scenarios of extremely low SNR, where the structural signal is nearly nonexistent, the phenomenon known as “Einstein from Noise” manifests in both 2D and 3D contexts. This phenomenon pertains to the inherent model bias within the reconstruction procedure, specifically in relation to the initial templates. In such cases, the reconstructed volume exhibits structural similarities to the initial template, even though the observations do not substantiate this outcome. The generation of a structured image from entirely noisy data has attracted considerable attention, particularly during a significant scientific debate regarding the structure of an HIV molecule [34, 19, 57, 52, 33]; for a comprehensive description and statistical analysis, see [3, 2]. Notably, our empirical evidence suggests that the “Einstien from Noise” phenomenon is more pronounced when adopting the MAP estimator compared to the MMSE estimator, implying that the MMSE approach is less vulnerable to the choice of the initial template. Furthermore, our experiment implies the advantage of using the Bayesian MMSE estimator over the MAP estimator is more significant in 3D structure reconstruction tasks compared to 2D image recovery, where the 3D setting is a problem of greater interest to researchers in structural biology.
@@ 5 Discussion and conclusions
In this work, we have introduced the Bayesian framework for enhancing orientation estimation for various applications in structural biology. The proposed approach offers greater flexibility and improved accuracy compared to existing methods, with the MMSE estimator as a prime example. This technique handles diverse structural conformations and arbitrary rotational distributions across sample sets. Our empirical results establish that the proposed MMSE estimator consistently surpasses the performance of the current methods, particularly in challenging low SNR environments, as well as when prior information on rotational distribution is available or approximately known. We provide a theoretical foundation to explain these performance gains. As rotation determination is crucial for both 2D and 3D reconstruction processes, we further illustrate how utilizing the MMSE estimator as a soft-assignment step in iterative refinement leads to significant improvements over hard-assignment methods. Moreover, the proposed Bayesian approach empirically offers enhanced resilience against the “Einstein from Noise” phenomenon, effectively reducing model bias and improving the overall reliability of structural reconstructions. Thus, our main recommendation is to adopt the Bayesian MMSE rotation estimator over the MAP estimator in related application scenarios. Already integrated into most software, the Bayesian rotation estimator can be easily implemented with minimal computational cost, offering improved accuracy and resilience for orientation determination and structure reconstruction.
@@
This study has primarily focused on the Bayes estimator with respect to the MSE loss function; however, exploring alternative loss functions beyond MSE presents an exciting avenue for future research. While the Bayes estimator with MSE has a closed-form solution, other loss functions typically lack analytic solutions, necessitating numerical estimation methods. Investigating the impact of different loss functions on 3D reconstruction accuracy would provide valuable insights. For instance, recent studies have demonstrated the effectiveness of the Wasserstein distance as a loss function for rotation estimation [50].
Another promising direction is the direct estimation of the rotational distribution from observations and integrating this information into the rotation estimation process or EM algorithms for 3D structure reconstruction [23]. Such an approach could significantly improve rotation estimation accuracy, particularly in low SNR regimes, as suggested by the results in Figure 2, and could potentially enhance volume reconstruction. Additional extensions could include modeling greater structural uncertainties, which is common in scenarios involving flexible proteins, and addressing general pose determination problems that encompass rotations and translations. Finally, leveraging the Bayesian framework to derive confidence regions for individual rotations offers another interesting research direction, potentially enhancing interpretability and reliability in rotation estimation tasks.
@@
The data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon request.
@@ Acknowledgment
T.B. is supported in part by BSF under Grant 2020159, in part by NSF-BSF under Grant 2019752, in part by ISF under Grant 1924/21, and in part by a grant from The Center for AI and Data Science at Tel Aviv University (TAD). We thank Eric Verbeke and Ruiyi Yang for their helpful discussions.
@@
Appendix A Proof of Proposition 2.1
By the definition in (2.11), assuming a uniform distribution of $\Lambda$ over $\mathsf{SO}(3)$ and that $\tilde{\Sigma}=\sigma^{2}I_{d\times d}$, we have:
$\hat{g}_{\mathsf{MMSE}}=$ $\mathbb{E}[g|y,\overline{V}]$
$=$ $\frac{\int_{\mathsf{SO}(3)}g\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\overline{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}\mathrm{d}g}{\int_{\mathsf{SO}(3)}\exp\Big{(}-\frac{1}{2}(y-\Pi(g^{-1}\cdot\overline{V}))^{\mathsf{T}}\tilde{\Sigma}^{-1}(y-\Pi(g^{-1}\cdot\overline{V}))\Big{)}\mathrm{d}g}$
$=\frac{\int_{\mathsf{SO}(3)}g\exp\Big{(}-\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}/2\sigma^{2}\Big{)}\mathrm{d}g}{\int_{\mathsf{SO}(3)}\exp\Big{(}-\left\|y-\Pi(g^{-1}\cdot\overline{V})\right\|^{2}/2\sigma^{2}\Big{)}\mathrm{d}g}.$ (A.1)
To prove the proposition, we apply the following known theorem [41, Theorem 5.10]:
Theorem A.1.

### v2

@@ 4.2 Empirical results for volume reconstruction and the “Einstein from Noise” phenomenon
A few observations can be made from Figures 4 and 5. First, in the case of high SNR (i.e., as $\sigma\to 0$), the volume reconstruction is similar whether using the MLE estimator or the MMSE estimator. This similarity is theoretically supported by Proposition 2.2. However, as the SNR decreases, the reconstructions diverge, with the volumes reconstructed using the MMSE estimator showing a better correlation with the true volume. Second, in scenarios of extremely low SNR, where the structural signal is nearly nonexistent, the phenomenon known as “Einstein from Noise” manifests in both 2D and 3D contexts. This phenomenon pertains to the inherent model bias within the reconstruction procedure, specifically in relation to the initial templates. In such cases, the reconstructed volume exhibits structural similarities to the initial template, even though the observations do not substantiate this outcome. The generation of a structured image from entirely noisy data has attracted considerable attention, particularly during a significant scientific debate regarding the structure of an HIV molecule mao2013molecular ; henderson2013avoiding ; van2013finding ; subramaniam2013structure ; mao2013reply ; for a comprehensive description and statistical analysis, see balanov2024einstein ; balanov2024confirmation ; balanov2025expectation ; balanov2025structure . Notably, our empirical evidence suggests that the “Einstein from Noise” phenomenon is more pronounced when adopting the MLE estimator compared to the MMSE estimator, implying that the MMSE approach is less vulnerable to the choice of the initial template. Furthermore, our experiment suggests that the advantage of using the Bayesian MMSE estimator over the MLE estimator is more significant in 3D structure reconstruction tasks compared to 2D image recovery, where the 3D setting is a problem of greater interest to researchers in structural biology.
5 Structural heterogeneity analysis
Understanding heterogeneity is central to revealing the dynamic behavior of macromolecular complexes in cryo-EM. Unlike traditional 3D classification methods that assume a small number of discrete states, modern approaches aim to capture continuous structural variability by embedding projection images into a latent conformational space. Here, we adopt the RECOVAR framework gilles2025cryo , which performs heterogeneity analysis using principal component analysis (PCA) based on a regularized estimate of the conformational covariance matrix. The key idea is to infer the covariance of 3D volumes directly from noisy 2D projection images, leveraging the fact that although the volumes themselves are unobserved, their projections contain sufficient statistical information.
5.1 Fixed-pose methods
A central computational assumption in RECOVAR is that the poses $\{\phi_{i}\}$ of all particle images are known and fixed, typically obtained from a consensus refinement procedure. This “fixed-pose” assumption underlies many modern heterogeneity analysis pipelines, including 3D Classification in cryoSPARC punjani2017cryosparc , 3D Variability Analysis (3DVA) punjani20213d , 3DFlex punjani20233dflex , cryoDRGN zhong2021cryodrgn , CryoDRGN-AI-fixed levy2024end , Opus-DSD luo2023opus , and RECOVAR gilles2025cryo . These methods are collectively referred to as fixed-pose methods, since they treat poses as known inputs during downstream inference. Among them, RECOVAR has recently emerged as one of the most effective approaches. According to the CryoBench evaluation jeon2024cryobench , which benchmarks multiple methods across diverse synthetic and experimental datasets, RECOVAR consistently achieves state-of-the-art performance in both structural resolution and latent space recovery. This motivates our choice to build upon the RECOVAR framework.
Despite its strong empirical performance, the fixed-pose assumption merits careful scrutiny. In practice, ground-truth poses are never directly accessible and cannot be recovered exactly, even if the underlying volume is known, due to noise and the intrinsic ill-posedness of the inverse problem. Most existing methods address this by using MLE pose estimates in real data. This gap between assumed and achievable pose accuracy motivates a central question: to what extent can more accurate pose estimation improve downstream heterogeneity recovery? In this work, we address this question by replacing MLE poses with our MMSE estimator within RECOVAR.
5.2 The mathematical model
The forward model (gilles2025cryo, , Eq. (1)) used in RECOVAR is expressed in the Fourier domain as
$\tilde{y}_{i}=C_{i}\hat{P}(\phi_{i})\tilde{V}_{i}+\tilde{\varepsilon}_{i},$ (5.1)
where $\tilde{y}_{i}\in\mathbb{C}^{N^{2}}$ is the observed Fourier-transformed image (with $N$ denoting the grid size along each spatial axis), $\tilde{V}_{i}\in\mathbb{C}^{N^{3}}$ is the (unknown) 3D Fourier volume corresponding to a particular conformation, $\hat{P}(\phi_{i})$ is the tomographic projection operator from 3D to 2D after a rigid-body motion $\phi_{i}=(g_{i},t_{i})$ with $g_{i}\in\mathrm{SO}(3)$ a rotation and $t_{i}\in\mathbb{R}^{2}$ an in-plane shift, $C_{i}$ is the contrast transfer function (CTF), and $\tilde{\varepsilon}_{i}$ is additive noise.
This formulation is a discretized and sampled version of our continuous forward model in (A.2). Specifically, the operator $\hat{P}(\phi_{i})$ acting on $\tilde{V}_{i}$ corresponds directly to $e^{-2\pi i\bm{\omega}\cdot t_{i}}\cdot(g_{i}\circ\tilde{V})(\omega_{1},\omega_{2},0)$, where each component reflects the same physical process, namely, in-plane shift, 3D rotation, and evaluation on the central slice in Fourier space.
To model structural variability, we treat each underlying volume $\tilde{V}_{i}$ as a random sample from a distribution over 3D conformations. We assume this distribution has a well-defined mean $\mu\in\mathbb{C}^{N^{3}}$ and covariance $\Sigma\in\mathbb{C}^{N^{3}\times N^{3}}$ given by
$\mu=\mathbb{E}[\tilde{V}],\quad\Sigma=\mathbb{E}[(\tilde{V}-\mu)(\tilde{V}-\mu)^{*}],$ (5.2)
where the expectation is taken over the conformational distribution. The objective of RECOVAR is to estimate both $\mu$ and $\Sigma$ from the observed projections $\tilde{y}_{i}$} with given poses $\phi_{i}$}, using regularized least-squares minimization (gilles2025cryo, , Eqns. (2)–(3)).
5.3 Numerical experiments
In our experiments, we consider a synthetic one-dimensional conformational transition, a standard benchmark for continuous heterogeneity analysis methods such as cryoDRGN zhong2021cryodrgn and RECOVAR gilles2025cryo . The corresponding ground-truth density maps along the conformational coordinate are shown in Figure 6(a). To generate the projection images, the underlying conformational states were uniformly sampled along the one-dimensional transition, and molecular orientations were sampled from a uniform distribution over $\mathrm{SO}(3)$. Figure 6(b) further displays a representative noisy projection under high-noise conditions.
To avoid storing and manipulating the full high-dimensional covariance matrix $\Sigma\in\mathbb{C}^{N^{3}\times N^{3}}$, RECOVAR first estimates a low-dimensional subspace of rank $r\ll N^{3}$ in which the covariance is approximated. Following (gilles2025cryo, , Appendix A.2), we estimate $\hat{\Sigma}$ via regularized least-squares, then form a rank-$d$ approximation $\hat{\Sigma}_{\mathrm{col}}$ by selecting a subset of columns using a greedy SNR-based criterion that ensures both high SNR and low inter-column correlation. An orthonormal basis $\tilde{U}\in\mathbb{C}^{N^{3}\times r}$ spanning $\hat{\Sigma}_{\mathrm{col}}$ is obtained via randomized SVD. The conformational mean $\hat{\mu}$ is also estimated at this stage. Each conformational volume is then represented as
$\tilde{V}_{i}=\tilde{U}z_{i}+\hat{\mu}$, (5.3)
where $\hat{\mu}$ is the estimated conformational mean from the first stage, and $z_{i}\in\mathbb{C}^{r}$ are low-dimensional coordinates. In RECOVAR, these coordinates are not explicitly recovered; their covariance matrix $\hat{\Sigma}_{\tilde{U}}\in\mathbb{C}^{r\times r}$ is estimated directly via a reduced least-squares problem (Eq. (14) in gilles2025cryo ), yielding a low-rank covariance approximation that characterizes the variability of $z_{i}$. The eigen-decomposition of $\hat{\Sigma}_{\tilde{U}}$ then provides the estimated principal components and eigenvalues in the reduced space. The complete procedure is detailed in (gilles2025cryo, , Algorithm 1).
We perform our analyses on a controlled synthetic dataset of 30,00030{,}000 projection images ($N=128$), simulated from a one-dimensional conformational transition discretized into $50$ equally spaced states to approximate continuous heterogeneity gilles2025cryo . We first evaluate the effect of orientation estimation on covariance recovery. The study is conducted in the high-noise regime of RECOVAR (Figure 6(b)), where even with ground-truth poses, the top $30$ principal components capture only about $40\%$ of the variance due to finite-sample and high-noise effects. We set $r=50$ and compare three pose sources: ground truth, MMSE estimates, and MLE estimates, keeping all other steps identical. For fixed-pose methods, poses are treated as given, so computational cost remains unchanged.
The accuracy of the estimated subspace is quantified by the percentage of total variance captured:
$\|U_{k}^{*}\Sigma U_{k}\|_{S}/\|\Sigma\|_{S},$
where $U_{k}$ contains the first $k$ estimated principal components, $\Sigma$ is the ground-truth covariance matrix, and $\Sigma^{1/2}$ is its matrix square root. Here, $\|A\|_{S}=\sum\sigma_{i}(A)$ is the Schatten $1$-norm of matrix $A$ with $\sigma_{i}$ denoting the singular values of $A$.
Figures 6(e) and 6(f) present the variance-capture results as described and the corresponding eigenvalue recovery, respectively. Figure 6(d) visualizes the first five recovered principal components. Overall, MMSE pose estimation yields principal components and eigenvalue spectra that are consistently closer to the ground truth than those obtained from MLE pose estimation.
Figure 6: MMSE pose estimation improves heterogeneity reconstruction over MLE. (a) Ground-truth density maps from a synthetic dataset simulating a one-dimensional conformational transition, a standard benchmark for methods such as cryoDRGN zhong2021cryodrgn and RECOVAR gilles2025cryo . Fifty equally spaced states approximate the continuous trajectory, with colors indicating positions along the pathway (only five states are shown in the figure for clarity). (b) Example of a noisy projection image from the high-noise regime used in this study, where even ground-truth poses recover only $\sim 40\%$ of the variance for the top 30 components due to finite-sample and high-noise effects. (c) Comparison of reconstructed structures using different pose priors (ground-truth, MMSE, MLE), showing MMSE yields structures closer to the true conformation state. (d) First five principal components estimated using ground-truth poses, MMSE pose estimates, and MLE pose estimates. MMSE results closely match the ground truth in structural detail, whereas MLE reconstructions degrade beyond the first modes, becoming blurrier and less representative of the underlying variability. (e) Subspace accuracy, measured as the percentage of total variance captured (as defined in (5.3)). Ground truth achieves $\sim 40\%$, MMSE $\sim 30\%$, and MLE $\sim 25\%$. (f) Eigenvalue recovery: both methods estimate the largest true eigenvalues well, but MMSE remains accurate for smaller eigenvalues, while MLE substantially overestimates them. These results show that MMSE pose estimation consistently yields more accurate recovery of conformational variability than MLE.
We further perform a more granular evaluation of structure reconstruction at the level of individual conformational states. For this analysis, we select all projection images corresponding to a specific conformational state and embed them into the latent conformational space ($z$-space, see (5.3)) using RECOVAR, conditioned on different pose priors: ground-truth, MMSE estimates, and MLE estimates. For each subset, the reconstruction corresponding to the mean embedding serves as an estimate of that specific conformational state. These reconstructions provide a direct measure of how different orientation estimates affect the accuracy of heterogeneity structure reconstruction. Figure 6(c) demonstrates that MMSE orientation estimates produce reconstructed structures closer to the ground-truth conformation than those obtained from MLE, further highlighting the advantage of MMSE pose estimation in high-noise heterogeneity analysis. We further compared the reconstructions obtained using MMSE and MLE orientation estimates against the ground-truth conformational state. The reported local FSC resolution scores (in Å) show that the 50% quantile resolutions are 11.45 Å and 11.55 Å for MMSE and MLE, respectively, while the 90% quantiles are 18.83 Å and 21.43 Å (also see gilles2025cryo Supplementary S.J.).
@@ 6 Discussion and conclusions
In this work, we have introduced the Bayesian framework for enhancing orientation estimation for various applications in structural biology. The proposed approach offers greater flexibility and improved accuracy compared to existing methods, with the MMSE estimator as a prime example. This technique handles diverse structural conformations and arbitrary rotation distributions across sample sets. Our empirical results establish that the proposed MMSE estimator consistently surpasses the performance of the current methods, particularly in challenging low SNR environments, as well as when prior information on rotation distribution is available or approximately known. We provide a theoretical foundation to explain these performance gains. As rotation determination is crucial for both 2D and 3D reconstruction processes, we further illustrate how utilizing the MMSE estimator as a soft-assignment step in iterative refinement leads to significant improvements over hard-assignment methods. Moreover, the proposed Bayesian approach empirically offers enhanced resilience against the “Einstein from Noise” phenomenon, effectively reducing model bias and improving the overall reliability of structural reconstructions. Thus, our main recommendation is to adopt the Bayesian MMSE rotation estimator over the MLE estimator in related application scenarios. Already integrated into most software, the Bayesian rotation estimator can be easily implemented with minimal computational cost, offering improved accuracy and resilience for orientation determination and structure reconstruction.
@@
The MMSE orientation estimator represents the most natural Bayes estimator in the context of cryo-EM, as it coincides with the M-step reconstruction of the EM algorithm. Nevertheless, alternative Bayes estimators associated with different loss functions could also be explored in future studies. A particularly promising direction is the direct estimation of the rotation distribution from observed images and the incorporation of this information into rotation estimation or EM-based 3D reconstruction algorithms janco2022accelerated ; xu2025misspecified . Such approaches could substantially improve rotation accuracy, especially in low-SNR regimes, as indicated by our results in Figure 2, and consequently enhance volume reconstruction quality. Further extensions may involve modeling structural uncertainties, particularly relevant for flexible or heterogeneous proteins, and addressing general pose estimation problems that jointly consider rotations and translations. Finally, leveraging the Bayesian framework to derive confidence regions for individual rotations presents an exciting opportunity to improve the interpretability and reliability of rotation estimation in cryo-EM.
@@
The detailed implementation and code are available at https://github.com/AmnonBa/bayesian-orientation-estimation.
@@ Acknowledgment
A.S. and S.X. are supported in part by AFOSR under Grant FA9550-23-1-0249, the Simons Foundation Math+X Investigator Award, NSF under Grant DMS 2510039, and NIH/NIGMS under Grant R01GM136780-01. T.B. is supported in part by BSF under Grant 2020159, in part by NSF-BSF under Grant 2024791, in part by ISF under Grant 1924/21, and in part by a grant from The Center for AI and Data Science at Tel Aviv University (TAD). We thank Marc Auréle Gilles for providing guidance on RECOVAR software. We also thank Marc Auréle Gilles, Eric Verbeke, and Ruiyi Yang for their helpful discussions.
@@
Appendix A Full cryo-EM model
The (homogeneous) cryo-EM imaging process can be described by a comprehensive generative model that incorporates the key physical components of image formation, including random orientations, in-plane shifts, tomographic projection, and contrast transfer function (CTF).
Formally, each observed 2D image $y_{i}$ is modeled as
$y_{i}(x_{1},x_{2})=h_{i}*T_{t_{i}}\Big(\Pi\big((g_{i}\circ V)(x_{1},x_{2},x_{3})\big)\Big)+\varepsilon_{i}(x_{1},x_{2}),~~~\text{for }(x_{1},x_{2},x_{3})\in\mathbb{R}^{3},$ (A.1)
where $V$ denotes the underlying 3D volume, $g_{i}$ denotes the random rotation, $\Pi$ is the tomographic projection operator (i.e., line integration along $x_{3}$ axis), $T_{t_{i}}$ denotes the in-plane shift operator with shift $t_{i}\in\mathbb{R}^{2}$, * denotes the 2D convolution in the image plane, $h_{i}$ is the point spread function (whose Fourier transform is commonly known as CTF), and $\varepsilon_{i}$ denotes the measurement noise. Equivalently, by applying the Fourier transform and using the Fourier Slice Theorem, we obtain the forward model in the Fourier domain:
$\tilde{y}_{i}(\omega_{1},\omega_{2})=\tilde{h}_{i}\cdot e^{-2\pi\mathbf{i}\bm{\omega}\cdot t_{i}}\cdot\left((g_{i}\circ\tilde{V})(\omega_{1},\omega_{2},0)\right)+\tilde{\varepsilon}_{i}(\omega_{1},\omega_{2}),~~~\text{for }\bm{\omega}=(\omega_{1},\omega_{2})\in\mathbb{R}^{2},$ (A.2)
where $\tilde{y}_{i},\tilde{h}_{i},\tilde{\varepsilon}_{i}$ are the 2D Fourier transforms of $y_{i},h_{i},\varepsilon_{i}$, respectively, and $\tilde{V}$ is the 3D Fourier transform of $V$.
In practice, both the images and the underlying volume are discretized and sampled on finite grids. Depending on whether the model is implemented in the spatial or Fourier domain, different but equivalent discretized versions arise. Specifically:
  • In the spatial domain, each image $y_{i}$ is represented as a vector in $\mathbb{R}^{N^{2}}$, and the 3D volume $V$ is discretized as an array in $\mathbb{R}^{N^{3}}$.
  • In the Fourier domain, we work with complex-valued data, where $\tilde{y}_{i}\in\mathbb{C}^{N^{2}}$ and $\tilde{V}\in\mathbb{C}^{N^{3}}$ are the discrete Fourier representations of the image and volume, respectively. Since the images and the volume are real-valued, it follows that $\tilde{y}_{i}$ and $\tilde{V}$ are conjugate symmetric, i.e., $\overline{\tilde{y}_{i}(\omega_{1},\omega_{2})}=\tilde{y}_{i}(-\omega_{1},-\omega_{2})$ and $\overline{\tilde{V}(\omega_{1},\omega_{2},\omega_{3})}=\tilde{V}(-\omega_{1},-\omega_{2},-\omega_{3})$.
Here, $N$ denotes the resolution of the reconstructed grid: each image consists of $N\times N$ pixels, and the volume is represented on an $N\times N\times N$ voxel grid.
In the heterogeneous case, the underlying volume $V$ may itself be a realization drawn from a distribution over conformational states $V_{i}\sim\mathcal{P}_{V}$ where $\mathcal{P}_{V}$ represents a (possibly low-dimensional) distribution capturing the variability in molecular conformations, such as discrete states, continuous manifolds. This setting gives rise to the heterogeneous cryo-EM problem, where the goal is to infer not just a single structure, but the distribution $\mathcal{P}_{V}$ from noisy 2D projections.
Appendix B Alignment between two noisy volumes
The model described in (2.1) can be generalized to address the alignment problem between two noisy volumes. In (2.1), a single observation $y$ was assumed to follow the model, with the volume $V$ considered deterministic and known. For aligning two noisy volumes, we now consider two observations $y_{i}$ for $i=1,2$, given by:
$y_{i}=g_{i}\circ V+\epsilon_{i},$ (B.1)
where $\epsilon_{i}\sim\mathcal{N}(0,\Sigma_{d\times d})$ for a positive-definite cocaraince matrix $\Sigma_{d\times d}$, and $g_{i}$ are drawn from a distribution $\Lambda$ over $\text{SO}(3)$. We assume that $\epsilon_{1},\epsilon_{2},g_{1},g_{2}$ are independent, and that the volume $V$ is deterministic but unknown. Furthermore, we assume that the noise statistics are invariant under the group actions of $\text{SO}(3)$, meaning the distributions of $\epsilon$ and $g\circ\epsilon$ are identical for any $g\in\text{SO}(3)$. The objective is to determine the “best” alignment $\hat{g}\in\text{SO}(3)$ between $y_{1}$ and $y_{2}$, according to criteria such as the MLE, MAP, or MMSE estimators.
From (B.1), the unknown volume $V$ can be expressed as:
$V=g_{1}^{-1}\circ(y_{1}-\epsilon_{1}).$ (B.2)
Substituting (B.2) into (B.1), yields:
$y_{2}$ $=g_{2}\circ V+\epsilon_{2}$ (B.3)
$=g_{2}\circ\left(g_{1}^{-1}\circ(y_{1}-\epsilon_{1})\right)+\epsilon_{2}$ (B.4)
$=(g_{2}\cdot g_{1}^{-1})\circ y_{1}-(g_{2}\cdot g_{1}^{-1})\circ\epsilon_{1}+\epsilon_{2}$ (B.5)
$=(g_{2}\cdot g_{1}^{-1})\circ y_{1}+\epsilon,$ (B.6)
where (B.5) follows from the associativity of group actions, and in(B.6), we have used $\epsilon\triangleq(g_{2}\cdot g_{1}^{-1})\circ\epsilon_{1}+\epsilon_{2}$. Since the noise statistics are invariant under group actions, it follows that $\epsilon\sim\mathcal{N}(0,2\Sigma_{d\times d})$.
Thus, the problem of estimating the relative rotation $g_{2}\cdot g_{1}^{-1}$ between $y_{1}$ and $y_{2}$ is similar to estimating $y_{2}$ under the assumption that $V$ is known. However, the covariance of the additive noise $\epsilon$ in this estimation is doubled compared to the case when $V$ is known.
Appendix C Proof of Proposition 2.1
By the definition of $\hat{g}_{\mathsf{MMSE}}$ (2.11), we have,
$\hat{g}_{\mathsf{MMSE}}=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\int_{\mathsf{SO(3)}}\left\|g-\hat{g}\right\|_{F}^{2}{\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g.$ (C.1)
Then, as $\left\|g\right\|_{F},\left\|\hat{g}\right\|_{F}$ are fixed for any $g,\hat{g}\in\mathsf{SO}(3)$, we have,
$\hat{g}_{\mathsf{MMSE}}$ $=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\int_{\mathsf{SO(3)}}\langle\hat{g},g\rangle{\mathbb{P}}_{\Lambda}(g|y,{V})\mathrm{d}g$ (C.2)
$=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\langle\hat{g},\hat{g}_{\mathsf{relax}}\rangle.$ (C.3)
As a consequence, (C.3) is equal to the rotation $\hat{g}\in\mathsf{SO}(3)$ that has the highest correlation with the relexed estimator $\hat{g}_{\mathsf{relax}}$ as in (2.12). It follows that
$\hat{g}_{\mathsf{MMSE}}=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,max}}\langle\hat{g},\hat{g}_{\mathsf{relax}}\rangle=\underset{\hat{g}\in\mathsf{SO}(3)}{\operatorname*{arg\,min}}\left\|\hat{g}-\hat{g}_{\mathsf{relax}}\right\|_{F}^{2}$ (C.4)
Thus, $\hat{g}_{\mathsf{MMSE}}$ is exactly the Procrustes procedure performed on the intermediate estimator $\hat{g}_{\mathsf{relax}}$.
Appendix D Proof of Proposition 2.2
We split the proof into two parts. First, we show that
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MAP}}=\hat{g}_{\mathsf{MLE}}.$ (D.1)
Second, we show that
$\lim_{\sigma\to 0}\hat{g}_{\mathsf{MMSE}}=\hat{g}_{\mathsf{MLE}}.$ (D.2)
For (D.1), the argument is straightforward. Since the prior density $\Lambda(g)$ is bounded below by some positive constant, the term $\log(\mathrm{d}\Lambda(g))$ becomes negligible as $\sigma\to 0$ compared to the other term. Consequently, the MAP estimator converges to the MLE in this limit.
For (D.2), we apply the following theorem [42, Corollary 5.11]:
Theorem D.1.